OpenStudy (anonymous):

Please help! Use the following to answer the question. Triangle a = 12, b = 15, c = 21. The area = a. 61 b. 88 c. 119 d. 301

OpenStudy (lgbasallote):

i suppose there's no figure?

OpenStudy (anonymous):

use hero formula here.

OpenStudy (anonymous):

@Igba so figure. @sami-21 What's the hero formula?

OpenStudy (anonymous):

\[\Huge Area=\sqrt{s(s-a)(s-b)(s-c)}\] where \[\Huge s=\frac{a+b+c}{2}\]

OpenStudy (anonymous):

can you find it now ?

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

you 're welcome :)